If ultraviolet light of wavelength 200. nm causes electrons to be ejected from a gold surface with a kinetic energy of 1.88 x 10^-19 J, what is the threshold frequency for ejection of an electron from gold? And what is the wavelength of the electrons?
1
Expert's answer
2019-10-17T07:31:42-0400

E=hv, v=E/h=1.88*10-19/6.626*10−34=2.8373076*1014 Hz

λ=c/v=299 792 458/2.8373076*1014=1.0566*10-6 m = 1056.6 nm